ABCD is a parallelogram
angle A= 2x+19, angle B= 3x+21. Find all angles

Answer by 303795(602)

A parallelogram has two pairs of congruent (same sized angles). The four angle sizes add up to 360 degrees.
Angle A and C will be the same size as will angles B and D.
angle A + angle B + angle C + angle D = 360
2x+19 + 3x+21 + 2x+19 + 3x+21 = 360
10x + 80 = 360
10x = 280
x=28
So angle A and C are 2*28+19 = 75 degrees
and angles B and D are 3*28+21 = 105 degrees.